# Flow Annealed Importance Sampling Bootstrap meets Differentiable Particle Physics

### v1

@@ 4 Results and Discussion
$\bm{\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}}$.
When training with the same number of target evaluations, rKLD and FAB with the replay buffer provide the best results with the highest importance sampling efficiency and a distribution that is most similar to the test data set (see Tab. 1). Additionally, the integral estimate computed with samples from the flow has the lowest standard deviations. When increasing the number of target evaluations, $\epsilon$ improves for rKLD more rapidly than for FAB w/o buffer. For FAB w/ buffer, samples have to be generated before the start of training to fill the replay buffer, which results in an offset in the number of target evaluations. During training, the normalizing flow parameters are updated by sampling from the prioritized replay buffer based on $\bar{w}_{\mathrm{AIS}}$, which aids the model significantly and the importance sampling efficiency improves at a faster rate compared to the flows trained with rKLD. Overall, flows trained with rKLD and both FAB approaches have no problem in modeling the complex 2D distribution, validated by histograms shown in Fig. 6.
@@
$\bm{e^{+}e^{-}\rightarrow t\bar{t},t\rightarrow W^{+}b,\bar{t}\rightarrow W^{-}\bar{b}}$.
@@ 5 Summary and Conclusion
We have transferred FAB [18], which utilizes AIS with HMC as a transition operator, from molecular configuration modeling to HEP, building on recently introduced, differentiable implementations of matrix elements. We have demonstrated that training FAB with a prioritized replay buffer is a promising approach for improving the efficiency for event generation, since passing the flow samples through an AIS chain guides training at early stages. In the future, we plan to scale this approach to more complex particle-interaction processes and assess the performance improvement in greater detail. Including information about the individual matrix element contributions via multi-channeling will likely lead to further performance improvements [9, 13, 14]. Additionally, our conceptual work needs to be combined with differentiable implementations of the parton density [36, 15] to be applicable to proton-proton collisions. Overall, this work has the potential to improve the quality and speed of sampling methods employed at the LHC and facilitate the efficient analysis of high-luminosity events.
@@
Acknowledgments and Disclosure of Funding
The authors thank the motivated reviewers for their valuable remarks and feedback which helped improve this paper. AK thanks Timothy Gebhard for extensive support in reviewing and formatting the draft and Nicole Hartman for proof-reading and correcting the manuscript. The authors thank Matthew Feickert for help with running MadGraph. AK thanks Peter Lepage for answering questions about running VEGAS+. The computational work described in this manuscript was performed on the RAVEN cluster of the Max Planck Computing and Data Facility (MPCDF). We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AK. MK is supported by the US Department of Energy (DOE) under grant DE-AC02-76SF00515. LH is supported by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094-390783311. We also thank the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311, as this work was partially performed at the MIAPbP workshop on Differentiable and Probabilistic Programming for Fundamental Physics.
Used software: This work has made use of many open-source Python packages, including blackjax [38], ComPWA [24], corner [39], distrax [40], haiku [41], JAX [29], FAB-JAX [18], MadJAX [17], MadGraph [42], matplotlib [43], numpy [44], pylhe [45] SymPy [28], TensorWaves [46], and VEGAS+ [34, 35]. The accessible color schemes in our figures are based on [47].
@@
Appendix A Schematic comparison of training methods
{subcaptionblock}
.5 Refer to caption {subcaptionblock}.5 Refer to caption
{subcaptionblock} Refer to caption
Figure 2: Training with fKLD
Figure 3: Training with rKLD
Figure 4: Training with FAB
Figure 5: Visualization of compared methods. The black arrows illustrate the forward pass, while the red arrows mark the backpropagation of the gradient to update the flow parameters $\theta$. For rKLD, the gradient propagation to the normalizing flow $q_{\theta}(x)$ (red) requires a differentiable target distribution $p(x)$ (blue). For FAB, the HMC updates compute the gradient of the distribution of interest, while the backpropagation of the gradient is stopped for the AIS samples $\bar{x}_{\mathrm{AIS}}$ and weights $\bar{w}_{\mathrm{AIS}}$. The dotted lines indicate that these steps do not occur when a gradient update is performed with samples from the replay buffer.
Appendix B Relationship of sampling efficiency and unweighting efficiency
@@
Importance sampling efficiency. The (importance) sampling efficiency $\epsilon$ (c.f. Eq. 2) can be derived from the effective sample size $ESS$ which allows the performance comparison of different Monte Carlo methods like Markov Chain Monte Carlo (MCMC) or importance sampling (IS) based on a set of weighted samples. If we draw $N$ samples from a less-than-ideal MCMC or IS proposal distribution $q(x)$, the $ESS$ indicates the number of independent samples that these would be equivalent to if drawn directly from the target distribution $p(x)$. Therefore, the effective sample size can be defined proportional to the ratio of the variance of an ideal MC estimator (i.e., sampling from the target $p$) and the variance of the less-than-ideal MCMC or IS estimator (i.e., sampling from the proposal $q$[32]. Through derivations outlined in [33], the $ESS$ can be related to the estimate
$\widehat{ESS}=N~{}\frac{\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\right)^{2}}{\frac{1}{N}\sum_{i=1}^{N}w_{i}^{2}}=\frac{\left(\sum_{i=1}^{N}w_{i}\right)^{2}}{\sum_{i=1}^{N}w_{i}^{2}}=\frac{1}{\sum_{i=1}^{N}\bar{w}_{i}^{2}}~{}.$ (4)
with the normalized importance weights $\bar{w}_{i}=w_{i}/\sum_{j=1}^{N}w_{j}$ [32]. The importance sampling efficiency $\epsilon$ as defined in Eq. 2 corresponds to the normalization of Eq. 4.
@@
Unweighting efficiency. The unweighting efficiency stems from the approach of refining samples obtained from the less-than-ideal proposal distribution $q(x)$ by keeping only a fraction of the samples in proportion to the ratio of the target distribution $p(x)$ and the proposal $q(x)$ [48]. The so-called raw weight $w_{i}=\frac{p(x_{i})}{q(x_{i})}$ corresponds to the importance weight in the MCMC and IS setting. The definition of the unweighting efficiency [48, 9, 49] as
@@
Relationship of efficiencies. In [32], different formulations of the generalized effective sample size are explored based on a set of required and desirable conditions. They conclude that defining the $ESS$ as $1/\sum_{i=1}^{N}\bar{w}_{i}^{2}$ (related to $\epsilon$) and $1/\max(\bar{w}_{1},...,\bar{w}_{N})$ (related to $\epsilon_{\mathrm{uw}}$) are proper and stable formulations. While both efficiencies suffer from large outlier weights, the unweighting efficiency is directly affected through $w_{\mathrm{max}}$ in the denominator as reported in [9, 10]. For this reason, we choose to report the performance based on the importance sampling efficiency in the main body of this work and provide additional estimates of the unweighting efficiency in Tab. 2. We do not apply bootstrap techniques that are designed to mitigate large outliers in the weight distribution [10].
@@
Appendix C Hyperparameters
Data generation and training.
[…]
Normalizing Flow.
We use normalizing flows based on coupling layers and rational-quadratic spline transformations [51] implemented in distrax [40] and haiku [41]. For the base distribution, we choose a uniform distribution over the unit hypercube. Since changes in the normalizing flow hyperparameters affect the expressivity of the density estimator equally for all investigated methods, we perform hyperparameter tuning only for rKLD training because it has the shortest run time. For $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$, we subsequently vary the number of spline bins $n_{\mathrm{b}}\in[4,6,8,10,12,14,16]$, the number of flow transformations $n_{\mathrm{t}}\in[4,6,8,10,12,14,16]$, and the number of neurons per hidden layer $n_{\mathrm{n}}\in[10,50,100,150,200,250,300]$ for the fully connected conditioner network with two layers. For the 8D ME, we compare loss values and efficiencies for $n_{\mathrm{b}},n_{\mathrm{t}}\in[6,8,10,12,14]$, and $n_{\mathrm{n}}\in[200,250,300,350,400]$. Considering the trade-off between expressivity and increase in optimization time, we select the following values based on the final training loss, validation loss, and efficiency: For $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$, the number of bins and transformations is $n_{\mathrm{t}}=n_{\mathrm{b}}=10$, and a conditioner network with two hidden layers and 100 neurons each is used. For the 8D ME, we set $n_{\mathrm{b}}=n_{\mathrm{t}}=14$, and use 400 neurons for each of the two hidden layers.
FAB.
Additional hyperparameters have to be selected for FAB. We use two (linearly spaced) intermediate distributions $M$ in the AIS sequence with a HMC transition operator containing a single iteration and three leap frog steps. The initial HMC acceptance rate is set to $p_{\mathrm{acc}}=0.65$ and is tuned dependent on the number of actually accepted samples. Furthermore, we have to specify the number of gradient updates per iteration $L=4$ (2D) and $L=2$ (8D) in the case of buffered training. The most important variable in hyperparameter optimization is the HMC step size, since it depends on the support of the target distribution. To obtain a suitable estimate, we start one FAB run with an arbitrary step size and observe how the value is adjusted during training. We adopt the converged values of $l_{\mathrm{init}}=0.05$ for the two-dimensional and $l_{\mathrm{init}}=0.005$ for the eight-dimensional ME for all subsequent runs.
Optimization.
Baseline
We compare our results with a physics-agnostic integral estimation method called VEGAS+ [34, 35]. This grid-based optimization method subdivides the support into a regular, rectangular grid and estimates the integral contribution of each subspace. With this information, the grid is iteratively updated to focus on regions with large contributions and the value for the integral estimate is calculated as a weighted average over multiple runs. Through a combination of stratified and importance sampling, VEGAS+ is fast and efficient and can account for correlations between dimensions. For the 2D ME, we choose a VEGAS+ grid with 64 bins in each dimension, a damping factor of $\alpha=0.5$, two warm-up iterations with $10^{3}$ evaluations per iteration, followed by 8 iterations with $2\cdot 10^{5}$ evaluations per iteration for the integral estimates. For the 8D ME, we doubled the grid to 128 bins per dimension, kept the damping factor, and increased the number of evaluations to $10^{4}$ for each of the two warm-up iterations. The integral estimate is obtained from 8 iterations with $10^{5}$ evaluations. We optimized these hyperparameters to the best of our knowledge. To make sure that VEGAS+ converged, we performed checks like increasing the number of evaluations by a factor of 10 per iteration which provided stable results and did not show significant deviations in the integral estimate for both examples. However, the low importance sampling efficiency and large deviation on the integral estimate for the 8D ME (c.f., Tab. 1) indicate that VEGAS+ struggles to adapt to the distribution. It is important to note that the results are obtained without multi-channeling since this physics information based on the contributing matrix elements is not provided to the normalizing flows either. Multi-channeling would lead to a performance improvement for both VEGAS+ as well as flow-based methods [13, 14]. Although the VEGAS+ optimization is significantly faster than training a normalizing flow, the flexibility of the latter results in higher importance sampling efficiencies. Since the main goal of training ME surrogates is to provide an optimally pre-trained model [14], the training time is amortized when generating a large number of events.
Appendix D Additional results
$\bm{\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}}$.
To illustrate the complexity of the two-dimensional ME, we show the target density evaluated on a grid as a Dalitz plot in Figure 6. A Dalitz plot is a physics-specific visualization of 2D amplitudes where the axes are chosen such that the histograms can be interpreted as an unnormalized density. Physics information can be extracted by comparing the Dalitz plot histograms based on simulated samples and measured data [26, 27]. As we train the normalizing flow on $[0,1]^{2}$, we apply a mapping to the flow samples to obtain the Dalitz plot representation. We provide histograms of $10^{6}$ unweighted samples obtained from the best performing normalizing flow for each investigated method to compare the sampling quality. This can qualitatively be compared to $\sim 2\cdot 10^{4}$ samples and their weights obtained from a converged VEGAS+ integrator. We observe that fKLD struggles to learn the horizontal and diagonal lines (“resonances”). Both normalizing flows trained with FAB produce histograms that exhibit the full structure of the ME and provide the best results. To additionally allow a quantitative visual comparison of higher-level quantities than the Dalitz plot, we include histograms of the invariant mass distribution for $m_{pK^{-}}^{2}$ and $m_{K^{-}\pi^{+}}^{2}$ in Fig. 7.
$\bm{e^{+}e^{-}\rightarrow t\bar{t},t\rightarrow W^{+}b,W^{-}\bar{b}}$.
We compare the corner plots of $10^{6}$ flow samples for each method to samples from the training data set generated with MadGraph [42] in Figure 8. The latter serve as a ground truth and illustrate challenging properties of MEs: Peaks at the boundary and correlations between dimensions are difficult for normalizing flows. Here, we visualize samples from the flow directly on the unit hypercube and one would need to apply the inverse RAMBO transformation [31] to obtain physical information for each outgoing particle. We can relate the dimensions of the unit hypercube to physical quantities that correspond to re-scaled versions of intermediate masses $\widetilde{\mathrm{M}}$, the cosine of azimuthal angles $\cos\widetilde{\mathrm{\theta}}$, and polar angles $\widetilde{\mathrm{\phi}}$. We observe that the normalizing flows especially deviate from the MadGraph distribution close to the phase space boundary which can be observed for all investigated methods. Correlations between dimensions, for example between the angles $\widetilde{\mathrm{\phi}}_{1}-\widetilde{\mathrm{\phi}}_{3}$ as well as $\widetilde{\mathrm{\phi}}_{2}-\widetilde{\mathrm{\phi}}_{3}$, appear to be challenging. While we report the performance metrics based on $\sim 10^{5}$ samples from an optimized VEGAS+ integrator in Tab. 1 and 2, we do not include the results in the corner plot since we observe large deviations explained by low efficiencies.

### v2

@@ 4 Results and Discussion
$\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$.
To illustrate the complexity of the two-dimensional ME, we show the target density evaluated on a grid as a Dalitz plot in figure 6. We provide histograms of $10^{6}$ unweighted samples obtained from the best performing normalizing flow for each investigated method to compare the sampling quality. This can qualitatively be compared to $\sim 2\cdot 10^{4}$ samples and their weights obtained from a converged VEGAS+ integrator. For fKLD, we observe that the horizontal line corresponding to the $K^{*}(892)$ resonance is not clearly visible compared to the results of VEGAS+, rKLD, and FAB. Furthermore, the crossing of the horizontal and diagonal lines appears as a blurred region for fKLD. This observation is consistent with figure 7, where we provide marginal histograms of the invariant mass distribution for $m_{pK^{-}}^{2}$ and $m_{K^{-}\pi^{+}}^{2}$. We average the results for three models trained with different random seeds and visualize the standard deviation per histogram bin. Overall, the binned samples obtained from the normalizing flows trained with fKLD show the largest standard deviations from samples obtained with rejection sampling. Flows trained with rKLD and FAB produce histograms that exhibit all features of the ME structure and provide the best results.
When training with the same number of target evaluations, rKLD and FAB with the replay buffer provide the best results with the highest importance sampling efficiency and a distribution that is most similar to the test data set (see table 1). Additionally, the integral estimate computed with samples from the flow has the lowest standard deviations. The change in the sampling efficiency when increasing the number of target evaluations is shown in figure 8(a). We can observe that $\epsilon$ improves for rKLD more rapidly than for FAB w/o buffer. All normalizing flow models show the same importance sampling efficiency of approximately $70\%$ for a low number of target evaluations. This corresponds to the efficiency obtained from rejection sampling with a uniform proposal distribution (shown as a black, dashed line in figure 8(a)). This behavior can be explained by the fact that the normalizing flows have a uniform base distribution and are initialized with an identity mapping, resulting in a uniform distribution at the beginning of training. For FAB w/ buffer, samples have to be generated before the start of training to fill the replay buffer, which results in an offset in the number of target evaluations. During training, the normalizing flow parameters are updated by sampling from the prioritized replay buffer based on $\bar{w}_{\mathrm{AIS}}$, which aids the model significantly and the importance sampling efficiency improves at a faster rate compared to the flows trained with rKLD. Since FAB evaluates the target several times in each iteration dependent on the chosen number of HMC steps and intermediate AIS distributions, it can benefit from additional information about the structure of the ME. Overall, flows trained with rKLD and both FAB approaches have no problem in modeling the complex 2D distribution.
@@
$e^{+}e^{-}\rightarrow t\bar{t},t\rightarrow W^{+}b,\bar{t}\rightarrow W^{-}\bar{b}$.
@@ 5 Summary and Conclusion
We have transferred FAB [23], which utilizes AIS with HMC as a transition operator, from molecular configuration modeling to HEP, building on recently introduced, differentiable implementations of matrix elements. We have demonstrated that training FAB with a prioritized replay buffer is a promising approach for improving the efficiency for event generation, since passing the flow samples through an AIS chain guides training at early stages. In the future, we plan to scale this approach to more complex particle-interaction processes and assess the performance improvement in greater detail. Including information about the individual matrix element contributions via multi-channeling will likely lead to further performance improvements [14, 17, 18]. Additionally, we plan to extend this conceptual work with differentiable implementations of the parton density [47, 22] to be applicable to proton-proton collisions. Overall, this work has the potential to improve the quality and speed of sampling methods employed at the LHC and facilitate the efficient analysis of high-luminosity events.
@@
The authors thank the reviewers of MLST and the NeurIPS workshop Machine Learning and the Physical Sciences, 2024 for their helpful suggestions and interesting questions which helped to improve the manuscript. AK thanks Timothy Gebhard for extensive support in reviewing and formatting the draft as well as Ludwig Burger and Nicole Hartman for proof-reading and correcting the manuscript. The authors thank Matthew Feickert for help with running MadGraph. AK thanks Peter Lepage for answering questions about running VEGAS+. The computational work described in this manuscript was performed on the RAVEN cluster of the Max Planck Computing and Data Facility (MPCDF). We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AK. MK is supported by the US Department of Energy (DOE) under grant DE-AC02-76SF00515. LH is supported by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094-390783311. We also thank the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311, as this work was partially performed at the MIAPbP workshop on Differentiable and Probabilistic Programming for Fundamental Physics. Used software: This work has made use of many open-source Python packages, including blackjax [53], ComPWA [29], corner [54], distrax [55], haiku [56], JAX [39], FAB-JAX [23], MadJAX [21], MadGraph [40], matplotlib [57], numpy [58], pylhe [59] SymPy [38], TensorWaves [60], and VEGAS+ [11, 49]. The accessible color schemes in our figures are based on [61].
@@
Appendix A Transformation of Dalitz plot to unit hypercube
While a Dalitz plot is the ideal visualization of the outgoing particles of a three-particle decay from a physics perspective, it is not ideal for normalizing flow training since the flow has to learn the support of the distribution. To prevent the normalizing flows from assigning non-zero probability density outside the phase space boundary, we map the Dalitz plot coordinates $(m_{pK^{-}}^{2},m_{K^{-}\pi^{+}}^{2})$ to the unit hypercube with a differentiable and invertible transformation. First, we convert the Dalitz plot into the square Dalitz plot by transforming one of the invariant masses, e.g., $m_{K^{-}\pi^{+}}^{2}$ to the helicity polar angle $\cos\theta_{K^{-}\pi^{+}}$ [37]. The mapping between the two variables is defined as
$\begin{array}[]{l}m_{K^{-}\pi^{+}}^{2}=m_{\pi^{+}}^{2}+m_{K^{-}}^{2}\\ \hskip 48.36958pt+~{}\frac{1}{2m_{pK^{-}}^{2}}\left(m_{\Lambda_{c}^{+}}^{2}-m_{pK^{-}}^{2}-m_{\pi^{+}}^{2}\right)\left(m_{pK^{-}}^{2}+m_{K^{-}}^{2}-m_{p}^{2}\right)\\ \hskip 48.36958pt-~{}\frac{1}{2m_{pK^{-}}^{2}}\cos\theta_{K^{-}\pi^{+}}\cdot\lambda^{\frac{1}{2}}(m_{\Lambda_{c}^{+}}^{2},m_{pK^{-}}^{2},m_{\pi^{+}}^{2})\cdot\lambda^{\frac{1}{2}}(m_{pK^{-}}^{2},m_{K^{-}}^{2},m_{p}^{2})~{}.\end{array}$
where the masses $m_{\pi^{+}},m_{K^{-}},$ and $m_{p}$ are constants and the kinematic Källén function is defined as
$\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz~{}.$
After applying the non-linear transformation from $(m_{pK^{-}}^{2},m_{K^{-}\pi^{+}}^{2})$ to $(m_{pK^{-}}^{2},\cos\theta_{K^{-}\pi^{+}})$, we can visualize the resulting distribution in a square Dalitz plot in figure 10. We can observe that the previously horizontal line gets mapped to an almost diagonal line.
Figure 10: Dalitz plot of $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$ decay and transformation to square Dalitz plot.
Appendix B Hyperparameters
Data generation and training.
[…]
Normalizing Flow.
We use normalizing flows based on coupling layers and rational-quadratic spline transformations [62] implemented in distrax [55] and haiku [56]. For the base distribution, we choose a uniform distribution over the unit hypercube. Since changes in the normalizing flow hyperparameters affect the expressivity of the density estimator equally for all investigated methods, we perform hyperparameter tuning only for rKLD training because it has the shortest run time. We only tune the most important hyperparameters related to the architecture of the normalizing flow since this is computationally costly. For $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$, we subsequently vary the number of spline bins $n_{\mathrm{b}}\in[4,6,8,10,12,14,16]$, the number of flow transformations $n_{\mathrm{t}}\in[4,6,8,10,12,14,16]$, and the number of neurons per hidden layer $n_{\mathrm{n}}\in[10,50,100,150,200,250,300]$ for the fully connected conditioner network with two layers. For the 8D ME, we compare loss values and efficiencies for $n_{\mathrm{b}},n_{\mathrm{t}}\in[6,8,10,12,14]$, and $n_{\mathrm{n}}\in[200,250,300,350,400]$. Considering the trade-off between expressivity and increase in optimization time, we select the following values based on the final training loss, validation loss, and efficiency: For $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$, the number of bins and transformations is $n_{\mathrm{t}}=n_{\mathrm{b}}=10$, and a conditioner network with two hidden layers and 100 neurons each is used. For the 8D ME, we set $n_{\mathrm{b}}=n_{\mathrm{t}}=14$, and use 400 neurons for each of the two hidden layers.
FAB.
Additional hyperparameters have to be selected for FAB. We use two (linearly spaced) intermediate distributions $M$ in the AIS sequence with a HMC transition operator containing a single iteration and three leap frog steps. The initial HMC acceptance rate is set to $p_{\mathrm{acc}}=0.65$ and is tuned dependent on the number of actually accepted samples. Furthermore, we have to specify the number of gradient updates per iteration $L=4$ (2D) and $L=2$ (8D) in the case of buffered training. The most important variable in hyperparameter optimization is the HMC step size, since it depends on the support of the target distribution. To obtain a suitable estimate, we start one FAB run with an arbitrary step size and observe how the value is adjusted during training. We adopt the converged values of $l_{\mathrm{init}}=0.05$ for the two-dimensional and $l_{\mathrm{init}}=0.005$ for the eight-dimensional ME for all subsequent runs. We do not optimize the size of the FAB replay buffer since we do not expect a significant influence on the performance. We only take into account that the replay buffer should be sufficiently large such that sampling data points from the buffer is informative for normalizing flow training. Therefore, we set the minimal size of the buffer to 10 times the batch size (i.e., $10^{4}$) and the maximal buffer size to 100 times the batch size (i.e., $10^{5}$).
Optimization.
Baseline.
We compare our results with the physics-agnostic integral estimation method VEGAS+ [11, 49]. This grid-based optimization method subdivides the support into a regular, rectangular grid and estimates the integral contribution of each subspace. With this information, the grid is iteratively updated to focus on regions with large contributions and the value for the integral estimate is calculated as a weighted average over multiple runs. Through a combination of stratified and importance sampling, VEGAS+ is fast and efficient and can account for correlations between dimensions. For the 2D ME, we choose a VEGAS+ grid with 64 bins in each dimension, a damping factor of $\alpha=0.5$, two warm-up iterations with $10^{3}$ evaluations per iteration, followed by 8 iterations with $2\cdot 10^{5}$ evaluations per iteration for the integral estimates. For the 8D ME, we double the grid to 128 bins per dimension, keep the damping factor, and increase the number of evaluations to $10^{4}$ for each of the two warm-up iterations. The integral estimate is obtained from 8 iterations with $10^{5}$ evaluations. We optimized these hyperparameters to the best of our knowledge. To make sure that VEGAS+ converged, we perform checks like increasing the number of evaluations by a factor of 10 per iteration which provide stable results and do not show significant deviations in the integral estimate for both examples. However, the low importance sampling efficiency and large deviation on the integral estimate for the 8D ME (c.f., table 1) indicate that VEGAS+ struggles to adapt to the distribution. It is important to note that the results are obtained without multi-channeling since this physics information based on the contributing matrix elements is not provided to the normalizing flows either. Multi-channeling would lead to a performance improvement for both VEGAS+ as well as flow-based methods [17, 18]. Although the VEGAS+ optimization is significantly faster than training a normalizing flow, the flexibility of the latter results in higher importance sampling efficiencies. Since the main goal of training ME surrogates is to provide an optimally pre-trained model [18], the training time is amortized when generating a large number of events.
Appendix C Relationship of sampling efficiency and unweighting efficiency
Importance sampling efficiency.
@@ Appendix C Relationship of sampling efficiency and unweighting efficiency
The (importance) sampling efficiency $\epsilon$ (c.f. 2) can be derived from the effective sample size $ESS$ which allows the performance comparison of different Monte Carlo methods like Markov Chain Monte Carlo (MCMC) or importance sampling (IS) based on a set of weighted samples. If we draw $N$ samples from a less-than-ideal MCMC or IS proposal distribution $q(x)$, the $ESS$ indicates the number of independent samples that these would be equivalent to if drawn directly from the target distribution $p(x)$. Therefore, the effective sample size can be defined proportional to the ratio of the variance of an ideal MC estimator (i.e., sampling from the target $p$) and the variance of the less-than-ideal MCMC or IS estimator (i.e., sampling from the proposal $q$[50]. Through derivations outlined in [51], the $ESS$ can be related to the variance of the importance weights via
$ESS=\frac{N}{1+\mathrm{Var}_{q(x)}[w]},$
and to the estimate
$\widehat{ESS}=N~{}\frac{\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\right)^{2}}{\frac{1}{N}\sum_{i=1}^{N}w_{i}^{2}}=\frac{1}{\sum_{i=1}^{N}\bar{w}_{i}^{2}}~{}.$ (4)
with the normalized importance weights
$\bar{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{N}w_{j}}~{}.$
The importance sampling efficiency $\epsilon$ as defined in 2 corresponds to the normalization of 4.
Unweighting efficiency.
@@
The unweighting efficiency stems from the approach of refining samples obtained from the less-than-ideal proposal distribution $q(x)$ by keeping only a fraction of the samples in proportion to the ratio of the target distribution $p(x)$ and the proposal $q(x)$ [64]. The so-called raw weight $w_{i}=\frac{p(x_{i})}{q(x_{i})}$ corresponds to the importance weight in the MCMC and IS setting. The definition of the unweighting efficiency [64, 14, 65] as
@@
Relationship of efficiencies.
In [50], different formulations of the generalized effective sample size are explored based on a set of required and desirable conditions. They conclude that defining the $ESS$ as $1/\sum_{i=1}^{N}\bar{w}_{i}^{2}$ (related to $\epsilon$) and $1/\max(\bar{w}_{1},...,\bar{w}_{N})$ (related to $\epsilon_{\mathrm{uw}}$) are proper and stable formulations. While both efficiencies suffer from large outlier weights, the unweighting efficiency is directly affected through $w_{\mathrm{max}}$ in the denominator as reported in [14, 16]. For this reason, we choose to report the performance based on the importance sampling efficiency in the main body of this work and provide additional estimates of the unweighting efficiency in table 2. We do not apply bootstrap techniques that are designed to mitigate large outliers in the weight distribution [16].